OpenStudy (anonymous):

Simplify the expression. –20 • 1/4 • 8 + 12 ÷ 4 A. –13 B. 13 C. –37 D. –7

OpenStudy (hyper):

Answer: C Solution: -20 x 1/4 x 8 + 12 /4 = (-5 x 8) + 3 = -40 + 4 = -37 = C